Calculate the electronegativity differences of the following and determine if they are largely ionic, non-polar covalent, or polar covalent?

S8, CaCl2, SOCl2, NaF, CBr4, BrF, LiF and AsH3
Answer #1

bona du) Concepts if E.No of diffrence > 2 800, it will be balay a) Sq structure. s - s-sos bond s-s S S SS Electro negativitMAA d) Not structuree Nauf bondo Na F. ko No difference of Na-f= h. N. off - 6. No of Na = 4-0.9=301 (32) 8 Soo bond is the lg) Ur estructuree lif. bond. Lif. BoNo diffrence of Li- f bond = F.Nooff - E.An of la = 4-123 so. bond as lonic. h) ds la str